Thomson Trucking has $21 billion in assets, and its tax rate is 25%. Its basic earning power (BEP) ratio is 10%, and its return on assets (ROA) is 6.25%. What is its times-interest-earned (TIE) ratio? Round your answer to two decimal places.
Answer #1

BEP = EBIT/Total Assets

0.10 = EBIT/21

EBIT = $2.1 billion

ROA = Net Income/Total Assets

0.0625 = Net Income/21

Net Income = $1.3125 billion

Net Income = (1 - tax rate)(EBIT - Interest Expense)

1.3125 = (1 - 0.25)(2.1 - Interest Expense)

Interest Expense = $0.35 million

TIE = EBIT/Interest Expense

TIE = 2.1/0.35

TIE = 6